OpenStudy (anonymous):

integrate coshx/(3sinx-2coshx) thank you in advance

OpenStudy (anonymous):

have u tried this urself yet?

OpenStudy (anonymous):

and is that sin suppose to be sin or sinh?

OpenStudy (anonymous):

yup, by definition, but im redoing it right now

OpenStudy (anonymous):

sorry sinh not sin

OpenStudy (anonymous):

this is a long integral

OpenStudy (anonymous):

i didnt get the answer, my final answer is \[\int\limits cotxdx\] which is wrong.

OpenStudy (anonymous):

yes it is, ok bear with me i ll go step by step

OpenStudy (anonymous):

let me know if u dont understand any step

OpenStudy (anonymous):

1.multiply numerator by denominator by -sech^3(x)\[\int\limits_{}^{} -\frac{ sech^2(x) }{ 2sech^2(x)-3sech^2(x)\tanh(x) }dx\]

OpenStudy (anonymous):

wait why -sech^3 x?

OpenStudy (anonymous):

it will make your integral less painful

OpenStudy (anonymous):

after this u can use u= tan(x) substitution

OpenStudy (anonymous):

how did u do ur first step?

OpenStudy (anonymous):

from my original answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i used it definition

OpenStudy (anonymous):

oh using definition will take too long

OpenStudy (anonymous):

just a question, where did sinh and cosh go?

OpenStudy (anonymous):

didnt u read my 1. step multiply num and denom by -sech^3(x)

OpenStudy (anonymous):

sorry a mistake in my calculation

OpenStudy (anonymous):

so should i go further with next step

OpenStudy (anonymous):

\[x+\frac{ 3 }{ 5 } \ln(\left| (e^{2x}-5)e^{-2x} \right|)+c\]

OpenStudy (anonymous):

by definition this should be ur answer^

OpenStudy (anonymous):

unfortunately it's not.. i

OpenStudy (anonymous):

i think i mess up in the middle but i already erased it, so i cant show you

OpenStudy (anonymous):

do u have to do it by definition, have u learnt substitution yet?

OpenStudy (anonymous):

yeah i have learnt subs but im not sure which one should be subs

OpenStudy (anonymous):

now continue after tanhx=u, the denom sech^2 x is change to 1-tanh^2 x?

OpenStudy (anonymous):

u want me to continue?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

2.let u=tanh(x) and sub this into eq. Remember sech^2(x)=1-tanh^2(x)\[\int\limits_{}^{}-\frac{ sech^2(x) }{ (3\tanh(x)-2)(\tanh^2(x)-1) }dx\]

OpenStudy (anonymous):

3.sub u=tanh(x) and du=sech^2(x)dx \[\int\limits_{}^{}-\frac{ 1 }{ (3u-2)(u^2-1) }du\]

OpenStudy (anonymous):

factor out constants in this case is -1 and use partial fraction \[-\int\limits_{}^{}(\frac{ 1 }{ 10(u+1)}-\frac{ 9 }{ 5(3u-2)}+\frac{ 1 }{ 2(u-1) })du\]

OpenStudy (anonymous):

r u getting it?

OpenStudy (anonymous):

wait, why the denom relationship change, i mean how can it became from ()-() to ()()

OpenStudy (anonymous):

i used partial fraction

OpenStudy (anonymous):

ohhh... ok i get it

OpenStudy (anonymous):

read the top part i am writing , i am also explaining what i did in that step

OpenStudy (anonymous):

then next, just integrate then subs the u, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you, n sorry im a slow learner

OpenStudy (anonymous):

its ok :)